OpenStudy (anonymous):

the ends of a base of an isosceles triangle are (-3,3) and (-4,4) and the length of each of the other sides is 13. find the coordinates of the third vertex

pooja195 (pooja195):

@zepdrix

pooja195 (pooja195):

?

zepdrix (zepdrix):

\[\large d_1=\sqrt{(x+3)^2+(y-3)^2}\]\[\large d_2=\sqrt{(x+4)^2+(y-4)^2}\]And we know that,\[\large d_1=d_2\]